Find the transfer function using the state and output equations:

(-10 1 |-27 0 1||x2|+ |0u; (-18 0 0l [x3. [X3]

Y = [1 0 0] |x2.
Answer #1

Trans fndion-- c [8T-A BD om dub A= BH -lo .B- - 27 1& C. 00] 7-A O 1 2 7 -18 O +27 + 8 Magnitudo of Csz A 1 -l +27 +18( 92-1-1) x0 (-1)| 278-C-18)+ 0 9+10) (9+10) (A) (278+18) firding Co factor mrateik of ma tix A (2) 1-+1 + 27 1 t27+18 1 &1+(1-0) 1 3+1 C31 -1 8+Io 3+2 - C8+10) C32=(Cリ +27 (8 tlo) 3+3 C33=) 8(8t10) +27 o8 +27) (22 factor Matri of Co 2 C13T C12 188A g0 [3I-A] 1A1 92 t103 9+10 t108 +278+18) 278+1o) Ho8+27 +18 -188 Lets Say X B0 <[8r-A TPe fs Calarlsethin ( e [31- 8- ,2NOW TPe C B 8+10 2+lo8+2 Ansear (93+1082 +278 +18)Please find the attached files to see the solution and for any confusion regarding the solution, comment and I will reply asap. Thank You.